y=5-4x
2x-3y=13
How to solve this equation

Answer by swincher4391(1107)

USE SUBSTITUTION METHOD
-----------------------
y = 5-4x ---> substitute this in for y in the second equation
2x-3(5-4x) = 13 ---> 2x-15+12x =13
14x = 28
x=2
---------
Plug x =2 back into either ORIGINAL equation
---------------
y = 5-4(2) ----> y = 5-8 = -3
y =-3
So our solutions are x=2, y=-3